OpenStudy (anonymous):

simplify, (-5)to the second power

OpenStudy (anonymous):

25

OpenStudy (anonymous):

\[(-a)^2 = a^2\] Here a = 5 Use this formula for getting the answer..

OpenStudy (anonymous):

Negative times negative is positive. So it is 5*5.

OpenStudy (anonymous):

@DanielHendrycks Please don't just give out the answer. That's not what the site is for.

OpenStudy (anonymous):

wired did you make this site lol way 2 go nice job lol

OpenStudy (anonymous):

@ChrisMaddy The Code of Conduct says this site isn't here just to give out answers: http://openstudy.com/code-of-conduct

OpenStudy (zzr0ck3r):

its common sence, unless you are just trying to get medals, and then I suggest you go outside more:)

OpenStudy (anonymous):

To a previous question of his, he didn't seem like he wanted to know why. http://openstudy.com/study#/updates/4ff0f787e4b03c0c488a0c9c

OpenStudy (anonymous):

Additionally, I often see people post the answer, then the explanation (since it could take longer and the person might go).

OpenStudy (anonymous):

What?

OpenStudy (zzr0ck3r):

reread?

OpenStudy (anonymous):

@DanielHendrycks It's usually best to lead someone to figure out the answer for themselves. Showing them how and why is never a bad idea, but if you just give out the answer first (and it's someone who's taking a test / quiz / doesn't want to do the work themselves), then it's not helping them to learn. You could always say "I know the answer, and here's how you can arrive at it" or something like that. @ChrisMaddy "What?" - Please explain what that is in reference to. Are you asking about your original question, or are you asking about the code of conduct on the site, or are you just asking what the answer is?